MATH 115: ELEMENTARY NUMBER THEORY HOMEWORK #7 WORKSHEET

JOHN VOIGHT

Cipher challenge! Below is a collection of three ciphertexts. Each one is encrypted in a different way. Each of the plaintexts reveals a hint for the next ciphertext. If you finish, fame is yours!

(1) HTWQW CSIHT WEMJX APLTA

(2) EIPTP LPUPV EADPN TQXCE

(3) e=39572139
n=59379373
begin 600 out
1437105
4244311
end

Date: August 3, 2004.